using the formula log, a = log a/log b find log^2 10000 round your answer to the hundreths place and show your work.
Answer by stanbon(75887)

using the formula log, a = log a/log b find log^2 10000 round your answer to the hundreths place and show your work.
---------
log(base 2)10000 = log(10000)/log(2) = 4/0.30103... = 13.29
-------
Cheers,
Stan H.